### v1

# DriveSOTIF: Advancing Perception SOTIF Through Multimodal Large Language Models

Abstract
Human drivers naturally possess the ability to perceive driving scenarios, predict potential hazards, and react instinctively due to their spatial and causal intelligence, which allows them to perceive, understand, predict, and interact with the 3D world both spatially and temporally. Autonomous vehicles, however, lack these capabilities, leading to challenges in effectively managing perception-related Safety of the Intended Functionality (SOTIF) risks, particularly in complex and unpredictable driving conditions. To address this gap, we propose an approach that fine-tunes multimodal language models (MLLMs) on a customized dataset specifically designed to capture perception-related SOTIF scenarios. Model benchmarking demonstrates that this tailored dataset enables the models to better understand and respond to these complex driving situations. Additionally, in real-world case studies, the proposed method correctly handles challenging scenarios that even human drivers may find difficult. Real-time performance tests further indicate the potential for the models to operate efficiently in live driving environments. This approach, along with the dataset generation pipeline, shows significant promise for improving the identification, cognition, prediction, and reaction to SOTIF-related risks in autonomous driving systems. The dataset and information are available: https://github.com/s95huang/DriveSOTIF.git
Index Terms:
SOTIF, LLM, VQA, LLM Agents, Multimodal LLM
I Introduction
In autonomous driving (AD), safety can be classified into functional safety and Safety of the Intended Functionality (SOTIF). Functional safety addresses issues with hardware and software related to a specific function, ensuring that the system operates correctly in response to its inputs. SOTIF, on the other hand, addresses safety concerns that arise not from system failures but from situations where the system performs as intended but may still lead to unsafe conditions due to external factors or system limitations [1].
Perception systems in autonomous vehicles (AVs), which are tasked with detecting, classifying, and predicting based on environmental stimuli, are particularly vulnerable to SOTIF-related challenges. Issues such as sensor limitations in adverse weather, difficulty in recognizing unusual road users, or interpreting rare and complex scenarios can lead to unsafe decisions [2]. The dynamic and unpredictable nature of real-world driving adds complexity, as sudden obstacles or erratic behavior from other road users can easily overwhelm current algorithms [3].
Detecting and addressing SOTIF risks are therefore essential for the safe operation of AVs. Existing methods adopt rule-based methods to quantify perception-related risks by measuring entropy levels and classifying them into specific categories [3, 4]. In contrast, human drivers can naturally assess the risk and uncertainty with certain objects when driving. This capability of assessing SOTIF-related risk and selecting the best action based on personal experience, knowledge, common sense, etc., remains absent from existing literature.
Recent advances in multimodal large language models (MLLMs) offer new opportunities to improve the understanding and management of SOTIF risks. By combining visual perception with language-based reasoning, these models can mimic the human ability to assess risk and uncertainty in driving scenarios.
II Literature Review
II-A Perception-related SOTIF
Perception SOTIF risks arise when an autonomous vehicle’s perception system functions correctly at the hardware and software levels but still fails to accurately interpret the driving environment due to external factors [3]. Since perception is fundamental for downstream decision-making and planning, any misinterpretation can propagate through the system, leading to unintended emergency braking, unsafe lane changes, or even collisions. These risks are further amplified by factors like sensor degradation due to rain, fog, snow, and glare, and uncommon object appearances [4]. Identifying and mitigating perception-related SOTIF risks is crucial for maintaining vehicle safety, particularly when an autonomous vehicle operates outside its predefined Operational Design Domain (ODD) [5, 6].
Current approaches to detect and mitigate perception-related SOTIF risks can be classified into rule-based and learning-based methods. Rule-based approaches, such as Fault Tree Analysis (FTA) [7], System Theoretic Process Analysis (STPA) [8, 9], and formal verification [10], attempt to identify perception-triggering conditions and system failures through structured models [9]. However, these methods struggle with dynamic and unpredictable environments, limiting their applicability to real-world driving conditions [11].
In contrast, learning-based and probabilistic uncertainty quantification techniques aim to detect and mitigate perception-related SOTIF risks through data-driven models. Methods such as Bayesian Risk Decomposition [12], deep ensemble [4, 3, 13], and sensor fusion techniques have been explored to quantify uncertainty in perception outputs and improve reliability [7]. Despite these advancements, learning-based methods often suffer from black-box behavior, making it challenging to interpret the underlying process, and poor generalization to edge cases, particularly in long-tail traffic scenarios.
II-B LLM in AD
Early research in the application of LLMs for autonomous driving primarily focused on enhancing interaction and decision-making (DM) through text-based LLMs. For instance, Cui et al. introduced the DriveLLM framework and demonstrated the feasibility of LLM-enhanced decision-making through experimental studies [14]. By integrating visual feature encoders with text-based LLMs, these models gain the ability to interpret visual inputs [15, 16, 17]. This combination allows the model to process and understand both visual and linguistic data, enabling more sophisticated reasoning and decision-making capabilities [18, 19].
In AD, this integration enhances the system’s capacity to interpret complex environments and respond to visual stimuli by combining visual inputs with language-based reasoning. Additionally, by combining image and textual information from other modules, such as vehicle states, safety, decision-making, planning, and control, the system can make more informed responses and better understand and adapt to dynamic driving scenarios [14, 20, 21]. This approach makes MLLMs well-suited for handling real-world complex driving scenarios, allowing autonomous systems to adjust to unpredictable conditions and improve safety and reliability [22, 23, 24, 25, 19, 20].
However, much of the existing research in the AD field relies on proprietary models, which require continuous access to internet. For open-source models, they are usually trained on general-purpose datasets like COCO Caption [26], VQAv2 [27], making them less suitable for AD domain applications.
One major issue with existing LLM-related autonomous driving datasets is the lack of diversity in driving scenarios. Datasets like DRAMA[28] and nuScenes-QA[23] are comprehensive in size but tend to focus on specific environments. Models trained on these datasets may face difficulties when geographical domain transfer or road conditions happen. Another significant limitation is the complexity and cost of annotations. Datasets like NuScenes-QA [23], Refer-KITTI [29] adopt a markup-based approach to generate template-based questions and answers. While this method balances well between dataset quality and cost, it results in a dataset that is not human-like. This limits the model’s ability to handle more open-ended or complex tasks that require a broader range of responses. Lastly, while datasets like DRAMA [28] emphasize safety-critical events, they don’t fully capture the complexity of real-world hazardous driving situations. In actual driving, safety risks come from unpredictable factors like sudden obstacles, changing weather, or erratic behavior from other road users. Current datasets often miss these scenarios, making it harder for models to perceive, understand, and respond to risks in real-time. Addressing this gap is essential for ensuring the safety and reliability of Level 3-5 autonomous vehicles.
III Overview
III-A Motivation and Contributions
Based on the literature review, despite the advancements in rule-based and learning-based approaches for detecting perception-related SOTIF risks, existing methods struggle with adaptability, interpretability, and robustness in dynamic driving environments. Rule-based techniques rely on pre-defined conditions that fail to generalize to unpredictable real-world scenarios, while deep learning-based methods often suffer from black-box behavior, poor generalization to edge cases, and vulnerability to adversarial conditions. Most importantly, these models lack contextual reasoning capabilities, making it difficult to understand, predict, and react appropriately to safety-critical situations. These limitations highlight the need for a novel approach that can detect and understand perception-related SOTIF risks.
Drawing inspiration from the way human drivers naturally perceive, reason, and act in complex environments, we propose leveraging MLLMs to address perception-related SOTIF risks. We fine-tune pre-trained MLLMs on DriveSOTIF, a domain-specific dataset carefully designed to capture the nuances of safety-critical driving scenarios. Through supervised fine-tuning (SFT), we enhance pre-trained MLLMs with domain-specific safety knowledge, allowing them to detect, reason, and respond effectively to safety-critical driving situations. This approach aims to enhance situational awareness, risk assessment, and real-time decision-making, leading to safer and more reliable autonomous vehicle systems. The key contributions of this study are summarized as follows:
  1. 1.
    Integration of MLLMs into SOTIF Research: We evaluate pre-trained MLLMs and enhance their performance through SFT on a custom domain-specific dataset. This comprehensive process—from dataset creation to model deployment—significantly improves the model’s ability to detect, reason about, and mitigate perception-related SOTIF risks in autonomous driving.
  2. 2.
    DriveSOTIF Dataset: The DriveSOTIF dataset is introduced as the first VQA and image captioning dataset specifically tailored for perception-related SOTIF risks. Its dataset generation pipeline can be applied to other autonomous driving tasks, making MLLM dataset creation more cost-effective and less prone to bias.
  3. 3.
    Deployment and Validation: Fine-tuned MLLMs were deployed and the effects of quantization and model size on real-time performance and compute requirements have been studied. Validation was conducted using data from the DriveSOTIF dataset, as well as real-world scenarios in Canada and China.
III-B Methodology
Figure 1: Supervised fine-tuneing of pre-trained multimodal LLMs on domain-specific dataset for enhanced perception SOTIF cognition and reaction
As illustrated in Figure 1, the proposed approach consists of three main stages:
  1. 1.
    Domain-Specific Data Curation (DriveSOTIF): We first constructed the DriveSOTIF datasetto to captures diverse, safety-critical driving scenarios, ranging from adverse weather conditions to rare object encounters. Each scenario is represented through images, visual-question answering (VQA) prompts, and image-captioning annotations that focus on perception-related SOTIF risks. The DriveSOTIF dataset is generated using a pipeline of collaborative MLLM agents, designed to generate the desired dataset efficievely with reduced hallucinatnions and bias. A round of human review is performed to ensure dataset quality and overall reliability.
  2. 2.
    Supervised Fine-Tuning: We begin with pre-trained MLLMs and fine-tune them on the DriveSOTIF dataset through a SFT procedure. We employ Low-Rank Adaptation (LoRA) and quantized LoRA (QLoRA), which allow us to a small portion of the model rather than retraining the entire model. This strategy significantly reduces computational overhead and memory usage, which is more practical for real-world applications.
  3. 3.
    Deployment and Validation: After fine-tuning, we benchmark the fine-tuned models on the test split of the DriveSOTIF dataset and compare their performance against the original pre-trained MLLMs. We then select two lightweight, high-performing models to examine how different model sizes and quantization levels affect real-time inference. Finally, we validate these models in real-world settings, including university campuses, city streets, and proprietary dashcam footage collected in China.
IV DriveSOTIF Dataset
IV-A Image Selection
For this research, we utilize the first batch of images from PeSOTIF [4], an object-detection dataset created to study the perception SOTIF problem in long-tail traffic scenarios. Specifically, the PeSOTIF dataset contains natural and artificial scenarios of perception SOTIF scenarios caused by environmental factors such as adverse weather conditions (e.g., rain, snow, fog), challenging lighting conditions (e.g., direct sunlight, glare from headlights), and object-related factors like unusual object appearances or unexpected object behaviors.
Today’s passenger vehicles often contain multiple cameras, posing a challenge for VQA tools that need to incorporate these inputs to effectively assist in BEV-based perception SOTIF problems. For a dataset newly introduced to SOTIF, we aim to start with simple one-image and one-question pairs. This approach allows for comparisons with existing VQA datasets and algorithms and encourages further research in this area.
TABLE I: DriveSOTIF Dataset Answer and Question Types
Type Description
Answer Types
Open-ended No fixed question structure and have multiple correct answers
Closed-ended Yes/no and other limited choices
Closed-ended Question Types
Uncertainty Uncertainty level with a certain object (Low, Medium, High)
Existence Does an object of class XXX exist?
Type What is type of object at the bottom left?
Count How many XXX are there?
Key object Is this object a key object?
For the VQA subtask, we designed our dataset such that it supports close-ended and open-ended Q&A, as shown in Table I. Closed-ended questions are divided into five categories: uncertainty, existence, type, counting, and key object. These questions typically use templates or examples for generation, providing straightforward answers with limited choices such as yes/no or other specific options.
Open-ended Q&A are designed for higher levels of reasoning, such as causal inference and complex reasoning. These questions do not follow a fixed structure and can have multiple correct answers, offering more flexibility and depth in responses. These questions are focused on SOTIF risk identification, explanation, and providing recommended actions.
To ensure a balanced representation of question types, we extend our dataset to include ten questions per image: five closed-ended and five open-ended.
This approach ensures a diverse set of questions and promotes a comprehensive evaluation of the dataset.
IV-B Dataset Generation through Collaborative LLM-Agents.
To automatically generate the VQA and image captioning datasets, we leverage a multi-agent framework utilizing proprietary LLMs with vision capabilities. In this study, we use OpenAI’s GPT-4v, GPT-4o, and Claude 3 Opus from Anthropic. The dataset generation pipeline, as illustrated in Figure 2, is divided into distinct roles for image captioning, question generation, answer generation, and validation/post-processing.
Figure 2: Proposed dataset generation pipeline through collaborative LLM Agents.
Image Captioning Agent: The image captioning agent aims to generate the caption for the image based on a prompt. Importantly, we alternates between two LLM backends for question generation and captioning. This approach is adopted to minimize bias that might arise from relying on a single proprietary LLM model. Sample images and captions are provided in Table II.
Question Generation Agent: The primary objective of this agent is to generate natural, human-like questions based on the input query, captions generated, along with 2D object detection and uncertainty information  [^1] Please refer to supplemental material for object detection and uncertainty quantification. . Specifically, the question-generating agent is instructed to generate questions with varying difficulties and types. For closed-end questions, the agent randomly selects the question type from Table I. For open-ended questions, the agent focuses on advanced QA like causal inference, including identifying and explaining the SOTIF risk and providing recommended action to prevent the risk, if possible. Sample questions generated are shown in Table II.
TABLE II: Sample labeled Questions & Answers and image captions from SOTIF-VQA dataset.
Task Image 1 Image 2
Image [Uncaptioned image] [Uncaptioned image]
Close-ended QA Q1: Is the visibility of road signs and pedestrians clear in this image? Q1: How many vehicles can you discern through the windshield?
A1: No A1: Four
Q2: Does the image show a clear view of the traffic lights? Q2: Is visibility reduced in this image due to weather?
A2: Yes A2: Yes
Open-ended QA Q: What specific environmental factor in the image degrades the perception ability of a driver or an autonomous driving system? Q: What are the perception-related SOTIF risks evident in this image due to rain?
A: The low sun angle causing lens flare and reduced visibility is the environmental factor degrading perception. A: The rain creates a wet and reflective road surface, reduced visibility, and potential glare from surrounding vehicle lights, which can degrade the ability of an automated driving system to accurately detect lane markings, other vehicles, and pedestrians.
Q: What technological improvement could help reduce the perception-related SOTIF risk due to direct sunlight? Q: What type of environment subset does this scenario belong to in the context of SOTIF problems?
A: Improving camera dynamic range and employing polarizing filters or advanced image processing algorithms to mitigate glare and enhance visibility could help. A: This scenario belongs to the ’rain’ subset within the environment category, which involves scenarios that degrade the perception ability of automated driving systems due to natural environmental factors.
Image Captioning A driver’s perspective of a busy intersection during sunset, highlighting the challenges of low sun glare in urban driving. Driving in heavy rain, the windshield is covered in water droplets, making visibility of the traffic ahead poor.
Answer Generation Agent: In the VQA dataset format, each question is answered by ten responses from ten annotators. While this approach ensures accuracy, it is both costly and time-consuming. To streamline this process, the answer generation agent is designed to provide ten answers for each question, given an image. To achieve this, we employ a model aggregation method that utilizes GPT-4V and Claude3-Opus. For each image, the image file and associated questions are fed into the LLMs. Instead of sequentially querying a single model multiple times, the system generate five answers for each question in a single pass through each model. This method is more efficient and cost-effective compared to sending the questions individually.
Validation and Post-processing Agent: As shown in Figure 2, newly generated answers are then sent to the validation and post-processing agent, which contains three post-processing mechanisms. The first and second mechanisms are to ensure that captions and questions generated are related to the image and perception SOTIF context. The third mechanism is to ensure that answers generated are correct based on the input image and question. If any inconsistencies are found, the agent regenerates the captions, questions or answers as needed. This validation process leverages OpenAI’s GPT4o as the LLM backend to ensure high-quality outputs that align with the perception SOTIF context.
IV-C Dataset Overview
The multi-agent system is implemented to generate the dataset using different temperature settings to ensure a variety of questions/answers and captions. After the dataset generation, the data entries are inspected to ensure correct structure and then organized and converted into the desired format. To ensure dataset quality and effectiveness, we conduct a round of human review and verification. This manual process identifies and corrects errors that automated systems might miss or generate.
Our dataset contains 1114 images, 1114 captions, 5570 questions, and 55700 answers  [^2] It should be noticed that collecting SOTIF-related scenarios is partially challenging due to their rarity and unpredictability in real-world settings, resulting in a relatively small dataset size. However, we are actively working with PeSOTIF authors to expand the dataset by adding new scenarios.. The image captioning dataset is organized into the COCO caption [26] format, while the VQA dataset is organized using VQAV2 [27] format [^3] For compatibility with recent research, we converted the image-captioning and VQA dataset into the LLaVA format. We also generated a Chinese version of the dataset by translating the original dataset using GPT-4o.. These datasets are split into train/validation and test sets at a ratio of 0.8 and 0.2, as shown in Table III.
TABLE III: Training and Validation split statistics.
Train/Valid Test
Number of questions 4455 1115
Number of answers 44550 11150
Number of captions 891 223
Average question length (words) 5.30 5.37
Average answer length (words) 3.51 3.44
Average caption length (words) 19.22 18.22
As shown in Table II, our generated dataset is highly related to perception-related SOTIF problem and and can be utilized to enhance MLLMs for VQA/image captioning tasks in AD. Furthermore, our dataset features an average caption length of 18.22 to 19.22 words, nearly double the average length of 10.5 words per caption in the COCO Caption [26] dataset. This increased length provides additional context and could potentially enhance model performance.
IV-D Evaluation Metrics
For evaluation, we primarily focus on two key aspects: the similarity between model-generated answers and human-generated answers, and the compatibility between the image and the model-generated response.
In image captioning, we used standard metrics such as BLEU, METEOR, ROUGE, CIDEr [30], and SPICE [31]. CIDEr calculates the average cosine similarity between candidate and reference sentences, while SPICE evaluates precision and recall based on scene graphs. Together, these metrics assess the accuracy, fluency, and relevance of the generated captions, providing a comprehensive evaluation of the model’s ability to describe visual scenes. For the VQA task, we relied on BLEU and ROUGE scores to measure how accurately and contextually the models answered visual questions. These metrics highlight the models’ proficiency in generating meaningful and relevant responses based on the given visual inputs.
V Model Benchmarking and Fine-tuning
V-A Benchmarking
For image captioning benchmarking, we evaluate the Blip [32] and Blip2 models [33] (pre-trained on the COCO Caption dataset [26]) across various model sizes. For VQA tasks, we benchmark a total of five open-sourced models, including LLaVA 1.5 [34], LLava-next[35], LLaVA-Video[36], Idelics2 [37] and newly released Qwen2-VL [38] that is developed based on Qwen2 [39].
These models are classified by architecture, model size, and quantization method, focusing on three specific variants: 2B, 7/8B, and 13/34B. The 2B variant is particularly suitable for mobile device deployment due to its lightweight and performance. The 7/8B model is well-suited for on-vehicle computing, where there is a need for both performance and efficiency. The 13/34B variant, due to its larger size and resource requirements, is more appropriate for cloud-based deployments. The model benchmarking and fine-tuning presented are completed on a desktop equipped with an Intel i9-10900K, 128GB of memory, and two Nvidia RTX 3090 GPUs.
V-B Model Fine-tuning
To enhance the performance of our benchmarked models, we adopt Parameter-Efficient Fine-Tuning (PEFT) techniques such as Low-Rank Adaptation (LoRA) [40] and Quantized LoRA (QLoRA) [41]. LoRA introduces low-rank matrices into each Transformer layer, allowing the model to adapt with far fewer trainable parameters while keeping the original model weights frozen. QLoRA builds on this by incorporating quantization, which further reduces memory requirements and enhances efficiency, all while maintaining performance levels. This makes it easier to adapt LLMs for specialized tasks without needing extensive hardware resources.
For the image captioning task, we fine-tune the Blip [32] and Blip2 [33] models using the LAVIS framework [42]. In the VQA tasks, we leveage LLaVA [34] and LLaMA-Factory [43] frameworks to fine-tune LLaVA 1.5 and Qwen2-VL family models due to their high performance scores when compared with other models.
To ensure the reproducibility of our results, we maintain consistent settings across fine-tuning and evaluation phases. As shown in Table IV, we use identical random seeds and hyperparameters, such as learning rate (LR), temperature, for all models.
TABLE IV: Settings for Fine-tuning and Evaluation
Parameter Value Parameter Value
Image Captioning
Max Epoch 5 LR Scheduler Linear Warmup Cos
Max Length 50 Initial LR 1e-5
Min Length 8 Minimum LR 0
Number of Beams 5 Warmup LR 1e-8
Image Size 364 Warmup Steps 1000
Freeze ViT True Weight Decay 0.05
VQA
Learning Rate 5e-5 Max New Tokens 512
Cutoff Length 1024 Temperature 0.95
Top-p 0.7 Quantization BitsandBytes
V-C Results and Analysis
V-C1 Qualitative Analysis
Qualitatively, we illustrate the results using three representative images from the test split, as depicted in Table V, where . These results include the labeled ground truth, the predictions generated by the baseline models, and those produced by the fine-tuned models. For image captioning, baseline model is BLip2 6.7B while fine-tuned Blip2 6.7B . For VQA, responses from raw and fine-tuned QWEN2VL-7B are presented.
From Figure V, it can be seen that baseline models for image captioning tend to give short, generic descriptions like “a car driving on a wet road” or “a car driving on a foggy road.” These captions, while accurate, often miss important details that add depth to the scene, such as specific road conditions or potential hazards. On the other hand, fine-tuned models offer much more detailed and context-aware descriptions. For example, instead of just noting the basic scene elements, they might say, “a driver’s view through a windshield on a rainy day, with red traffic lights ahead,” drawing attention to factors like visibility and potential obstructions. In foggy conditions, they might highlight “low visibility”, offering a clearer understanding of the challenges faced by the driver.
In the VQA tasks, the baseline models generally perform adequately when answering simple, close-ended questions but tend to struggle with more complex, open-ended queries. In contrast, the fine-tuned models demonstrate a significant improvement in handling open-ended questions by providing more detailed and context-aware answers.
For instance, in response to a question about rain’s impact on perception, the fine-tuned model identifies challenges like “reflections and distortions caused by the rain on the windshield,” offering a deeper understanding of the driving environment.
In more complex scenarios, such as an overturned vehicle, the fine-tuned model provides more thorough answers, highlighting issues like blocked lines of sight and shadow zones that could complicate the vehicle’s response. These detailed insights showcase the fine-tuned model’s superior ability to handle complex, safety-critical situations.
TABLE V: Performance comparison of baseline models (Base), fine-tuned models (FT), and ground truth (GT) for visual question answering (VQA) and image captioning (IC) tasks.
Task Type Image 1 Image 2 Image 3
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IC Base: a view of a car driving on a wet road a car driving on a foggy road a car that is upside down on the side of the road
FT: a driver’s view through a windshield on a rainy day, with red traffic lights visible on the road ahead a driver’s view through a car’s windshield on a foggy road, with low visibility and limited visibility a car flips over on a busy road, causing traffic congestion and hindering visibility for drivers on a busy city street
GT: View from a vehicle’s dashboard camera showing a rainy and blurred roadway scene with visible windscreen wiper, a truck ahead, and blurry tail lights Intense fog conditions on a road pose a challenge for drivers, with only faint outlines of vehicles ahead visible An overturned car in the middle of a busy street highlights potential hazards in urban traffic conditions
VQA Q1: Is there precipitation affecting visibility in the scenario depicted? Is the visibility reduced in this driving scenario due to fog or mist? Is the overturned vehicle a key object in the scene?
Base: Yes Yes Yes
FT: Yes Yes Yes
GT: Yes Yes Yes
Q2: What weather condition is depicted in the image that could affect perception systems in vehicles? What environmental factor is degrading the perception ability in this image? What perception-related SOTIF risk is present due to the vehicle’s unusual position in the scene?
Base: rainy weather, which can affect the perception systems by reducing visibility, making it more difficult for the systems to accurately detect and track objects. The image shows that the environmental factor degrading perception ability is fog camera’s field of view may not capture the full extent of the scene
FT: The image shows heavy rain conditions, which can significantly degrade the performance of perception systems in vehicles by scattering and reflecting light The perception ability is degraded by heavy fog, which significantly reduces visibility on the road. The overturned vehicle poses a risk as it is not commonly expected to be in that position on the road, which can lead to misclassification or delayed recognition by autonomous driving systems.
GT: The image shows heavy rain conditions which can degrade the visibility and accuracy of vehicle perception systems by scattering light and obscuring road markings and other vehicles. The environmental factor shown in the image is fog, which significantly reduces visibility and therefore degrades the perception ability of drivers and automotive sensors. The overturned vehicle presents a high SOTIF risk by unexpectedly altering the driving environment, which could lead to inappropriate responses from an autonomous driving system unprepared for such irregular situations.
Q3: What specific aspect of the environment in this image can autonomously driven vehicles struggle to interpret correctly? What manual safety precaution should be taken by drivers in foggy conditions like in the image? How could the presence of other vehicles near the overturned car influence the autonomous vehicle’s response?
Base: The image shows a road with rain, which can make it difficult for autonomous vehicles to interpret the environment correctly. Rain can cause the road surface to become slippery, making it more difficult for vehicles to maintain traction and control. Additionally, rain can reduce visibility, making it harder for vehicles to see obstacles and other vehicles on the road. Decrease speed to allow for better reaction time. Increase following distance to give enough time to react if the vehicle in front stops suddenly. Use headlights to improve visibility, but avoid high beams as they can blind other drivers. Use fog lights if available for additional illumination. Keep windshield wipers off to avoid blurring visibility. Ensure mirrors are clear and properly adjusted. Avoid sudden braking, acceleration, or turning. Use turn signals to inform other drivers of your intentions. Stay in your current lane and avoid lane changes unless absolutely necessary. The presence of other vehicles near the overturned car could influence the autonomous vehicle’s response in several ways. First, the autonomous vehicle may need to adjust its speed and direction to avoid colliding with the overturned car and other vehicles on the road. Second, the autonomous vehicle may need to communicate with other vehicles on the road to ensure that they are aware of the situation and take appropriate actions. Finally, the autonomous vehicle may need to coordinate with emergency services to ensure that they are aware of the situation and can respond accordingly.
FT: The reflections and distortions caused by the rain on the windshield can lead to misinterpretation of the road condition and the position of other vehicles or pedestrians, which are critical for safe navigation Drivers should reduce their speed, increase following distances, use fog lights if available, and ensure their vehicle’s headlights and windshield are clean for better visibility. Other vehicles could block line-of-sight or create shadow zones, complicating the vehicle’s ability to assess the situation and determining the appropriate response.
GT: Autonomously driven vehicles can struggle to correctly interpret lane markings and other road signs obscured by water and reflections on the wet road surface, increasing the risk of navigation errors. Drivers should reduce their speed, use fog lights if available, maintain a safe following distance, and stay focused on the road to adjust quickly to potentially unseen hazards. Other vehicles close to the overturned car might limit maneuvering options for an autonomous vehicle and complicate the scene understanding, necessitating advanced decision-making algorithms that incorporate predictions of other drivers’ behaviors.
V-C2 Quantitative Analysis
The quantitative results, shown in Tables V-C2 and V-C2, demonstrate that fine-tuned models consistently outperform their baseline counterparts across all evaluation metrics.
TABLE VI: Image Captioning Scores
Method Size Trainable Parameters (M) BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr SPICE
Baselines
Blip [32] Small 223.97 0.348 0.219 0.132 0.082 0.134 0.315 0.138 0.086
Large 446.29 0.349 0.227 0.139 0.083 0.137 0.324 0.140 0.091
[…]
6.7B 107 0.398 0.270 0.169 0.106 0.146 0.338 0.176 0.096
Fine-Tuned Models
Blip [32] Small 223.97 0.659 0.464 0.318 0.217 0.208 0.414 0.411 0.143
Large 446.29 0.671 0.463 0.312 0.206 0.210 0.399 0.410 0.145
[…]
6.7B 107 0.723 0.509 0.353 0.248 0.225 0.421 0.536 0.155
TABLE VII: VQA Scores Comparison: Baseline vs. Fine-tuned Models
Method Size Quan. BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
Base FT Base FT Base FT Base FT
LLaVA 1.5 [34] 7B - 18.95 46.06 26.09 40.40 5.70 13.14 14.80 30.86
13B 8 bit 18.83 46.50 26.21 41.27 5.69 14.47 14.96 31.85
[…]
13B 8 bit 12.59 - 21.72 - 4.71 - 10.58 -
34B 4 bit 13.72 - 22.75 - 5.07 - 11.74 -
[…]
7B - 25.55 46.36 29.38 40.63 7.14 13.27 18.50 30.80
In image captioning, the fine-tuned Blip2 6.7B model obtained the best results, achieving ROUGE-L, CiDEr, and SPICE scores of 0.421, 0.536, and 0.155, respectively. This marks an increase of 24.56%, 204.54%, and 61.46% compared to the baseline version. Similar improvements were observed in METEOR, ROUGE, CIDEr, and SPICE scores, indicating the models’ enhanced ability to generate more accurate and contextually relevant descriptions of complex scenes. Notably, the fine-tuned Blip2 2.7B model achieved comparable performance, highlighting the importance of balancing model size with computational efficiency and performance needs.
For the VQA task, among baseline models, the newly released Qwen2-7B achieved the highest scores in BLEU, and ROUGE metrics. Additionally, it can be seen that larger models generally achieve better performance. For example, the 34B version of LLaVA-Next, implemented using 4-bit precision, outperformed its 13B and 7B variants. A similar trend is observed when comparing the Qwen2-VL and LLaVA 1.5 models. For the fine-tuned models, the LLaVA 1.5 - 13B model deployed with 8-bit precision achieved the highest scores across all metrics, which is expected given its larger size. Interestingly, the Qwen2-VL 2B model delivered similar level of performance with only 1/7 the number of parameters, making it a highly efficient option for deployment.
Similar to image captioning, fine-tuning also led to significant performance gains in VQA tasks. The BLEU-4 score for the LLaVA 1.5 model 13B version increased by 146.95% from 18.83 to 46.50, significantly improving the model’s accuracy and contextual understanding in handling open-ended questions. Similar improvements can be seen across all other metrics, showing that the fine-tuned models consistently generate more relevant and contextually appropriate responses.
V-C3 Discussion
Based on the results from both the quantitative and qualitative analyses, it is evident that fine-tuning leads to significant improvements in model performance across image captioning and VQA tasks. Fine-tuned models consistently generate more accurate, detailed, and contextually relevant outputs compared to their baseline counterparts.
Additionally, our findings demonstrate that the DriveSOTIF dataset, despite its relatively small size of 1,114 images, can effectively enhance the performance of MLLMs through SFT. This highlights the potential of using focused, domain-specific datasets to significantly improve model performance. Lastly, the proposed dataset generation approach can be adopted in other autonomous driving tasks, offering an efficient pathway for adopting MLLMs into specialized applications.
VI Experiments
VI-A Model Deployment and Testing
As illustrated in Figure 3, common deployment framework for MLLM models include Huggingface, vLLM [44], and TensorRT-based methods [45]. These models can be deployed locally on devices, in the cloud, or through multi-access edge computing (MEC) or edge computing setups. At a high level, the deployment process can be simplified to an API request/response service, as shown in Figure 3. In this setup, input images and text are sent to the API, where deployed MLLM processes them and returns the response.
In this research, the fine-tuned versions of Qwen2-VL 2B and 7B models were chosen for deployment due to their efficient performance and compact size. These models were deployed using the Lamma-Factory framework [43] through Huggingface, with a batch size set to 1. Both models are merged with their corresponding fine-tuned adaptive weights and deployed on a single RTX 4090 GPU.
For testing, 111 images from the test split of DriveSOTIF dataset were used. Each image was paired with the following input prompt: “We are studying perception-related SOTIF problems for autonomous driving vehicles. What are the SOTIF risks of this image, and what are the reasons and suggested actions? Provide an informative, natural, human-like, and accurate response. Do not make up information.”
Figure 3: Overview of multi-model LLM deployment.
Model Mem. [GB] Quant. Mean [s] Median [s] Std Dev [s] Min [s] Max [s] 2B 8.0 - 0.82 0.80 0.11 0.59 1.23 6.3 8bit 6.45 6.22 0.85 4.68 8.80 4.8 4bit 2.21 2.17 0.36 1.44 3.30 7B 19.5 - 2.78 2.33 1.18 0.98 6.02 13.2 8bit 10.04 8.44 4.31 5.00 23.42 9.2 4bit 3.57 3.09 1.55 1.44 7.98
TABLE VIII: Processing time statistics for 2B and 7B models.
Figure 4: Processing time per image
The peak GPU memory usage and processing time per image are recorded and plotted in Figure 4.
The statistical analysis is performed, and the result is presented in Table VIII. Based on our results, quantization primarily reduces GPU memory usage but does not improve real-time performance. This is reasonable since only the pre-trained LLM is quantized, while other computations still operate with higher precision. As a result, for the Qwen2-VL 2B model, GPU memory usage decreases from 8.07 GB (full precision) to 6.33 GB (8-bit) and further down to 4.78 GB (4-bit). A similar trend can be observed in the 7B model.
In terms of real-time performance, for the Qwen2-VL 2B model, the non-quantized version had an average processing time of 0.82 seconds per image, with a median of 0.80 seconds and a standard deviation of 0.11 seconds. When the model was quantized to 8-bit precision, the processing time increased to 6.45 seconds per image. In contrast, using 4-bit precision reduces it to 2.21s per image. A similar pattern can be seen in the Qwen2-VL 7B variant.
The significant slowdown for 8-bit occurs because the consumer-grade GPU we used isn’t optimized to handle calculations with lower-bit-width tensors directly. As a result, the system has to perform extra steps to convert these lower-bit tensors back into float values before performing the necessary computations [46]. These additional steps introduce significant overhead, dramatically slowing down the process. In contrast, 4-bit quantization is slightly faster than 8-bit. This happens as 8-bit often requires intermediate conversions, while 4-bit tensors can be packed more efficiently, leading to reduced memory bandwidth overhead and improved performance. Nonetheless, with dedicated hardware and further optimization and support of the deployment pipeline, this slowdown could be mitigated.
VI-B Real World Data Evaluation
For real-world evaluation, as shown in Figure VI-B, we assessed the fine-tuned models using data collected in Waterloo, Ontario, Canada, along with proprietary dash-cam footage from China. This evaluation is conducted through a ROS code that queries the MLLM API with random questions from a list, along with image data from collected data in the form of ROSBag or video.
In detail, we used image and video data from the University of Waterloo’s WATonoBus project [47] and the Canadian Adverse Driving Conditions (CADC) dataset [48]. The WATonoBus is an autonomous shuttle operating along a 2.7-kilometer, two-lane University Ring Road that encircles the campus [14]. The CADC dataset was gathered by an autonomous vehicle platform navigating various areas around Waterloo.
The first scenario features a road maintenance crew working on a snowy road inside the university campus with a snowplow in operation. The snowplow, being a rare object, along with the workers and traffic cones, pose challenges for perception systems. The second scenario involves a truck and trailer preparing to merge into the ego vehicle’s lane on a snowy, slippery residential road, where visibility is reduced due to the adverse weather conditions. These two scenarios effectively demonstrate the SOTIF challenges faced by autonomous vehicles in both public roads and university campus environments.
In China, we evaluated the models on two challenging scenarios. The first scenario involved a child hidden in a flattened box lying on the ground, with only a small portion of his foot visible. This scenario is particularly difficult, as the hazard is subtle and easily missed by both human drivers and autonomous systems. In this case, the human driver successfully stopped in time, avoiding an accident. Our aim is to examine whether the fine-tuned MLLM can detect these tiny yet critical visual cues to prevent such accidents. The second scenario took place at night, where the ego vehicle was driving on a wet, reflective road surface due to rain. The glare from the headlights of an oncoming vehicle further complicated the situation, making it difficult to perceive the road and surrounding environment clearly. In this scenario, an electric bicycle crossed the road median unexpectedly and collided with the ego vehicle, resulting in an accident.
TABLE IX: Sample labeled Questions & Answers and image captions from SOTIF-VQA dataset.
Waterloo - Snow Plowing, University of Waterloo campus Waterloo - Merging Scenario, city of Waterloo
[Uncaptioned image] [Uncaptioned image]
[…]
A: The risks include potential misidentification of road boundaries due to snow, reduced contrast between the road and its surroundings, and the presence of workers and equipment which could be obscured by snow or not well-illuminated. A: The perception-related SOTIF risk in this image is caused by heavy snowfall, which can reduce visibility and obscure road markings, signs, and other vehicles.
China - Kid hiding in a box China - Strong Vehicle Front Lights, with reflective road
[Uncaptioned image] [Uncaptioned image]
[…]
A: The SOTIF risks include the potential for an autonomous vehicle’s sensors to misinterpret the presence and position of objects on the road, such as the boxes and the bicycle, due to the cluttered and irregular street setup. This could lead to inappropriate vehicle responses or failures to detect critical road users. A: The perception SOTIF risks include failure to detect lane markings, nearby vehicles, or pedestrians due to reduced visibility. The recommended action includes activating hazard lights and slowing down, and possibly engaging in manual driving if the perception system fails.
VI-C Results and Discussion
As can be seen in Table VI-B, in our experiments, fine-tuned MLLMs performed effectively in challenging environments, yet it also highlighted critical areas for enhancement. Notably, in Waterloo’s snowy conditions, the MLLM successfully managed complex scenarios with obscured road markers and temporary obstacles like snowplows and maintenance crews. This performance highlights the model’s robust feature recognition capabilities under adverse weather conditions, proving its potential in dynamic real-world applications.
In the Strong Vehicle Front Lights scenario, despite intense glare from oncoming vehicle headlights, the MLLM successfully identified potential risks and provided reasonable recommended actions. The model recognized that the glare could significantly reduce visibility, increasing the likelihood of missing nearby vehicles or pedestrians. In response, it suggested actions that could directly address the visibility issue and enhance safety, such as slowing down, increasing following distance, or activating hazard lights.
This demonstrates the model’s capability to not only detect risks but also understand the causal links between challenging conditions and safety threats, providing actions that could effectively reduce risks in real-world driving situations. With only 8 GB of GPU memory, the fine-tuned Qwen2-VL 2B model can be deployed to vehicles and replace the API-based vision model for local, (near) real-time visual risk assessment. This eliminates privacy concerns, cybersecurity risks, and reliance on internet access.
Pairing this model with a decision-making framework like DriveLLM[14] enables autonomous systems to factor SOTIF risks into real-time decision-making. This integration improves safety in complex environments and potentially prevents fatal accidents like the child-hidden-in-the-box scenario, where conventional perception systems may fail to anticipate hidden risks. Most importantly, deploying a fine-tuned model that supports continuous improvement significantly enhances adaptability. By continuously updating the dataset and applying techniques like SFT and Kahneman-Tversky Optimization (KTO)[49], the MLLM can progressively refine its performance and extend to other tasks such as scenario understanding and ODD assessment.
Conversely, the MLLM’s performance in detecting subtle hazards, such as the partially visible child in a box during the China experiments, revealed some limitations. While the model detected unusual patterns associated with the obscured visual cues, it did not accurately capture and interpret the partially or fully obscured child. This limitation is noteworthy as it highlights a scenario where even humans might potentially miss. One possible reason is the current capacity of the visual encoders used in the MLLM.
While increasing encoder size could improve detection, it would also result in a larger model and slower real-time performance. A more efficient approach could be to use a lightweight visual encoder combined with a compact language model. By integrating temporal information from multiple frames, the model could better track subtle changes and movements, improving hazard detection without significantly increasing computational load.
VII Discussion
VII-A Domain Transfer
From the labeled response presented in Table II, proprietary LLM models demonstrate the ability to provide accurate and relevant answers to complex, domain-specific questions related to AD perception and SOTIF risks, even without prior specialized training. This highlights their effectiveness in zero-shot and one-shot scenarios, where little or no task-specific data is available.
Moreover, as seen in Tables V and VI-B, SFT processes enable baseline models to incorporate domain-specific knowledge. These fine-tuned models perform reasonably well when answering questions about SOTIF challenges for vehicles operating in both Waterloo, Canada, and China. This indicates that the models can effectively transfer learned knowledge across different environmental contexts, adapting to diverse traffic scenarios, weather conditions, and regulatory landscapes.
That said, there are certain limitations to this approach. While domain transfer allows for generalization across different contexts, model performance may be less consistent in environments with extreme conditions or significantly different operational parameters. Although the models adapt well to many contexts, some scenarios may still require additional fine-tuning to maintain accuracy and reliability, especially when facing highly specific regulatory or environmental challenges.
VII-B Hallucinations
During the dataset generation process, we encountered cases of hallucinations in the labeled responses obtained before the validation agent. For instance, when asked whether a specific object was present in an image, half of the ten generated responses indicated “yes,” while the other half said “no.” This inconsistency highlights a common issue with large vision-language models (LVLMs), known as object hallucination [50].
After applying post-processing with LVLMs for validation, these hallucinations, particularly fact-conflicting hallucinations [50], can still appear in the dataset. Fine-tuned models may still inherit these hallucinations from their base LLM models.
Addressing hallucinations is a complex and open problem, especially in applications where responses must be both precise and trustworthy. Recent advancements, such as Fine-grained Direct Preference Optimization (FDPO) [51], Visual Contrastive Decoding (VCD) [52] and methods like HallE-Switch [53] have shown promise in reducing hallucination rates by 40-55%. Despite these improvements, hallucinations remain a persistent challenge. In the context of SOTIF, several strategies can be employed to mitigate these issues. These include integrating rule-based systems, incorporating human-in-the-loop validation, and developing hybrid models that combine statistical learning with formal verification methods.
VII-C Future of LLM in SOTIF
As current research on the application of LLMs in the SOTIF domain remains limited, this section lists potential future uses of both text-based and multimodal LLMs.
First, a universal description system for SOTIF scenarios can be developed to provide standardized descriptions of SOTIF scenarios. Building on this, a custom-trained or fine-tuned text-based LLM can be integrated to generate JSON-based representations for scenario detection, communication, and risk identification.
Text-based and multimodal LLMs could also be employed to efficiently generate SOTIF test cases, following systematic and structured testing plans. By leveraging LLM capabilities, these test plans can be progressively refined so that failure in one test case informs subsequent testing, ultimately reducing the total number of tests required.
Another promising application is the generation of highly targeted test cases. Given a specific SOTIF risk, LLMs can focus on creating scenarios that probe edge cases and rare, uncommon conditions, ensuring that the most challenging and potentially dangerous situations are thoroughly evaluated.
Multimodal LLMs can assist in evaluating rule-based or learning-based SOTIF mechanisms. By providing advanced scene understanding and cognitive abilities, MLLMs can refine manually defined SOTIF evaluation algorithms and crosscheck the results from conventional systems, ensuring that SOTIF risks are properly assessed from multiple angles. Additionally, MLLMs can select specific scenarios of interest based on input prompts.
The potential application of multimodal LLMs lies in integrating data from multiple sensor types, such as lidar, radar, and bird-eye-view (BEV) images. By encoding these inputs into feature embeddings and mapping them alongside textual prompts, LLMs could directly assess SOTIF-related risks at the feature level. This would provide a more detailed understanding of risks, eliminating the reliance on processed detection results and potentially identifying hazards earlier in the pipeline.
VIII Conclusion and Future Work
Our research demonstrates the feasibility of applying SFT techniques to MLLMs using the DriveSOTIF dataset to enhance perception-related SOTIF risk detection in autonomous driving systems. By introducing a small domain-specific dataset intended for model fine-tuning, we were able to evaluate strengths and limitations of MLLMs in perception-related SOTIF problems. Benchmark results indicate that our fine-tuned MLLMs outperform baseline models both qualitatively and quantitatively.
These improvements lead to enhanced risk identification and situational awareness, even under challenging conditions such as adverse weather and occlusion-heavy scenarios, as observed in data collected from both China and Canada. Real-world performance testing further confirms that the deployed models achieve an average inference time of 0.82 seconds per image, making them suitable for low- to moderate-speed applications like urban navigation, automated parking, and traffic jam assistance.
More importantly, by integrating fine-tuned MLLMs with LLM-based decision-making frameworks, autonomous vehicles can incorporate perception risks into real-time operations, ultimately improving road safety. Moreover, the ability of MLLMs to interpret complex scenarios through natural language reasoning paves the way for more transparent and interpretable safety assessments. Lastly, the on-device or MEC-based deployment of these lightweight models can lead to a privacy-preserving, onboard-cloud hybrid LLM deployment for next-generation autonomous vehicles.
For future work, we plan to expand the DriveSOTIF dataset by incorporating additional modalities such as BEV images, LiDAR point clouds, and radar data, which will further enhance the spatial-temporal reasoning capabilities of MLLMs. Our study also highlights several challenges that warrant further investigation, including effectively managing occlusion-heavy scenarios, scaling the technology for high-speed and safety-critical applications, and reducing the dependence on human review while ensuring data quality and minimizing hallucinations. Addressing these challenges will require collaborative efforts at the intersection of natural language processing and autonomous driving research.
Acknowledgment
The authors would like to acknowledge the financial support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and the grant provided by OpenAI.