### v2

# DriveSOTIF: Advancing SOTIF Through Multimodal Large Language Models

Abstract
Human drivers possess spatial and causal intelligence, enabling them to perceive driving scenarios, anticipate hazards, and react to dynamic environments. In contrast, autonomous vehicles lack these abilities, making it challenging to manage perception-related Safety of the Intended Functionality (SOTIF) risks, especially under complex or unpredictable driving conditions. To address this gap, we propose fine-tuning multimodal large language models (MLLMs) on a customized dataset specifically designed to capture perception-related SOTIF scenarios. Benchmarking results show that fine-tuned MLLMs achieve an 11.8% improvement in close-ended VQA accuracy and a 12.0% increase in open-ended VQA scores compared to baseline models, while maintaining real-time performance with a 0.59-second average inference time per image. We validate our approach through real-world case studies in Canada and China, where fine-tuned models correctly identify safety risks that challenge even experienced human drivers. This work represents the first application of domain-specific MLLM fine-tuning for SOTIF domain in autonomous driving. The dataset and related resources are available at github.com/s95huang/DriveSOTIF.git
Index Terms:
SOTIF, LLM, VQA, LLM Agents, Multimodal LLM
I Introduction
In autonomous driving (AD), safety is commonly classified into functional safety and Safety of the Intended Functionality (SOTIF). Functional safety concerns failures in hardware or software that result in unsafe operation. In contrast, SOTIF addresses hazards that occur not due to malfunctions, but when the system operates as intended yet produces unsafe outcomes because of external factors or inherent limitations [1].
Perception systems in autonomous vehicles (AVs),which are responsible for detecting, classifying, and predicting based on environmental inputs, are especially vulnerable to SOTIF-related risks. Factors such as sensor limitations in adverse weather, difficulty recognizing unusual road users, or rare and complex scenarios can result in unsafe decisions [2]. The dynamic, unpredictable nature of real-world driving further complicates matters: sudden obstacles or erratic behavior by other road users can easily overwhelm current algorithms [3].
Identifying and mitigating SOTIF risks is therefore essential for safe AV operation. Current methods primarily use rule-based approaches to quantify perception-related risks, for example by measuring output entropy and classifying risks into predefined categories [4]. Learning-based methods, on the other hand, suffer from black-box behavior, poor generalization to edge cases, and vulnerability to adversarial conditions. In contrast, human drivers intuitively assess risk and uncertainty for specific objects as they drive, drawing on personal experience, knowledge, and common sense—capabilities that remain absent from existing automated systems.
Recent advances in multimodal large language models (MLLMs) provide new tools to address these cognitive gaps. MLLMs offer three distinct advantages over traditional approaches: (1) open-world cognition that leverages pre-trained knowledge to interpret novel scenarios; (2) causal reasoning that traces how environmental factors cascade into safety risks; and (3) contextual inference that interprets social dynamics and behavioral patterns beyond conventional algorithmic reach [5, 6]. For example, an MLLM can infer that sun glare may obscure lane markings, recognize that people gathering with umbrellas likely indicates a bus stop, or understand how rain, construction, and nighttime conditions interact to create heightened safety risks. Crucially, MLLMs provide explainable risk reasoning via natural language, articulating why scenarios pose specific risks and recommending actionable responses—capabilities essential for safety validation and user trust.
TABLE I: Perception-Related SOTIF Datasets
Dataset Year Size Image LiDAR Text Description
PeSOTIF[7] 2022 1126 frames 2D object detection dataset for long-tail traffic scenarios with diverse weather conditions and uncommon object appearances.
[…]
Motivated by human drivers’ cognitive abilities, we propose a novel framework that leverages MLLMs to address perception-related SOTIF risks. Our approach fine-tunes pre-trained MLLMs on DriveSOTIF, a domain-specific dataset carefully designed to capture safety-critical driving scenarios. Through supervised fine-tuning (SFT), we enhance these models with specialized knowledge that allows them to detect, reason about, and effectively respond to safety-critical driving situations. This approach directly improves situational awareness, risk assessment, and decision-making for safer and more reliable autonomous vehicles. The key contributions of this study are summarized as follows:
  1. 1.
    Novel Integration of MLLMs into SOTIF Research: We introduce a comprehensive approach that integrates pre-trained MLLMs with supervised fine-tuning (SFT) on a tailored, domain-specific dataset. Our method enhances the capabilities of these models to accurately detect, reason about, and effectively mitigate perception-related SOTIF risks.
  2. 2.
    DriveSOTIF Dataset: We introduce DriveSOTIF, the first visual question answering and image captioning dataset for perception-related SOTIF challenges. Our novel multi-agent generation pipeline reduces annotation costs while maintaining high quality, providing a scalable approach for developing MLLM datasets for autonomous driving.
  3. 3.
    Deployment and Validation: Fine-tuned MLLMs were deployed and the effects of quantization and model size on real-time performance and compute requirements have been studied. Validation was conducted using data from the DriveSOTIF dataset, as well as real-world scenarios in Canada and China.
II Literature Review
II-A Perception-Related SOTIF
Perception-related SOTIF risks occur when an autonomous vehicle’s perception system functions correctly at both the hardware and software levels yet fails to accurately interpret the environment due to external factors [3]. Because perception underpins downstream tasks such as decision-making and planning, errors at this stage can propagate, leading to hazardous outcomes including false emergency braking, unsafe lane changes, or collisions. These challenges are particularly acute in adverse conditions—rain, snow, fog, glare—or in the presence of rare or unusual objects [7], especially when vehicles operate outside their defined Operational Design Domain (ODD) [11, 12].
Mitigation approaches can be broadly categorized as rule-based or learning-based. Rule-based methods—such as Fault Tree Analysis (FTA) [4], System-Theoretic Process Analysis (STPA) [13, 14], and formal verification [15]—use structured frameworks to model how specific failures can lead to unsafe behaviors. While effective for high-level system modeling, these methods often fail to capture the unpredictable and dynamic aspects of real-world environments [16].
Learning-based and probabilistic techniques focus on modeling perception uncertainty using data-driven methods. Bayesian Risk Decomposition [17], deep ensembles [7, 3, 18], and sensor fusion have shown promise in quantifying and mitigating uncertainty in perception outputs [4]. However, these models typically function as black boxes, limiting interpretability and making it difficult to diagnose failure modes. Their performance also tends to degrade in rare or edge-case scenarios that fall outside typical training distributions.
II-B MLLMs in Autonomous Driving and Dataset Challenges
Multimodal large language models (MLLMs) are attracting increasing attention in autonomous driving due to their ability to jointly process visual and textual information [19]. This enables the integration of visual inputs—such as camera feeds—with map data, vehicle state, and textual cues, leading to a more holistic understanding of the driving environment [20, 21].
From a SOTIF perspective, MLLMs are especially promising for addressing perception limitations. Their cross-modal reasoning allows them to compensate when one sensor stream is degraded—for example, when camera images are obscured by glare or LiDAR signals are weakened by fog. By integrating complementary sensory inputs, MLLMs can infer scene context, improving situational awareness and reducing the risk of misinterpretation under challenging conditions [22, 21, 23].
Figure 1: Overview of the proposed methodology for supervised fine-tuning of pre-trained MLLMs on DriveSOTIF dataset.
However, the effective deployment of MLLMs in SOTIF applications is currently hampered by dataset limitations. Most available SOTIF-related datasets, summarized in Table I, are focused on object detection tasks. For example, PeSOTIF [7] and CODA [8] target 2D detection in rare and long-tail scenarios, while SOTIF-PCOD [9] provides synthetic LiDAR point clouds. Of these, only CODA-LM [10] includes textual annotations for vision-language models, but its primary focus is not direct SOTIF risk assessment.
In addition, existing MLLM-focused driving datasets present further challenges. Datasets like DRAMA [24] and NuScenes-QA [25], while large and well-annotated, are geographically or scenically constrained, which limits their ability to generalize to new environments. Many also rely on template-based question-answer generation [25, 26], resulting in interactions that lack the nuance and diversity of natural human queries. This restricts the models’ capacity for open-ended reasoning and complex situational judgment, which are essential for safety-critical autonomous driving applications.
III DriveSOTIF
We introduce DriveSOTIF, a domain-specific dataset designed to capture perception-related SOTIF scenarios for autonomous driving. As illustrated in Figure 1, our approach combines dataset generation with supervised fine-tuning of pre-trained MLLMs.
III-A Image Selection
For this research, we utilize the first batch of images from PeSOTIF [7], an object-detection dataset created to study the perception SOTIF problem in long-tail traffic scenarios. Specifically, the PeSOTIF dataset contains natural and artificial scenarios of perception SOTIF scenarios caused by environmental factors such as adverse weather conditions (e.g., rain, snow, fog), challenging lighting conditions (e.g., direct sunlight, glare from headlights), and object-related factors like unusual object appearances or unexpected object behaviors.
Today’s passenger vehicles often contain multiple cameras, posing a challenge for VQA tools that need to incorporate these inputs to effectively assist in BEV-based perception SOTIF problems. For a dataset newly introduced to SOTIF, we begin with simple image-question pairs. This approach allows for comparisons with existing VQA datasets and algorithms and encourages further research in this area.
TABLE II: DriveSOTIF Dataset Answer and Question Types
Type Description
Answer Types
Open-ended No fixed question structure
Closed-ended Yes/No and other limited choices
Closed-ended Question Types
Uncertainty Uncertainty level with a certain object (Low, Med, High)
Existence Does an object of class (class) exist?
Type What is the type of (object) at the bottom left?
Counting How many (objects) are there?
Key object Is this (object) a key object?
Figure 2: Proposed dataset generation pipeline through collaborative LLM Agents.
For the VQA subtask, we designed our dataset such that it supports close-ended and open-ended Q&A, as shown in Table II. Closed-ended questions are divided into five categories: uncertainty, existence, type, counting, and key object. These questions typically use templates or examples for generation, providing straightforward answers with limited choices such as yes/no or other specific options.
Open-ended Q&A target higher-level reasoning, including causal inference and scenario explanation. These questions do not follow a fixed structure and can have multiple correct answers, offering more flexibility and depth in responses. These questions are focused on SOTIF risk identification, explanation, and providing recommended actions.
To ensure a balanced representation of question types, we include five questions per image, with 2–3 closed-ended and 2–3 open-ended. This approach ensures a diverse set of questions and promotes a comprehensive evaluation of the dataset.
III-B Dataset Generation through Collaborative LLM-Agents.
To automatically generate the VQA and image captioning datasets, we leverage a multi-agent framework using vision-capable large language models (LLMs), specifically OpenAI’s GPT-4v, GPT-4o, and Anthropic’s Claude 3 Opus. The dataset generation pipeline, illustrated in Figure 2, includes distinct roles for image captioning, question generation, answer generation, and validation/post-processing. All prompts used in the dataset generation and validation process are provided in S.III of the supplemental material.
Image Captioning Agent: The image captioning agent aims to generate captions based on input prompts. Importantly, we alternate between two LLM backends for question generation and captioning. This approach is adopted to minimize bias that might arise from relying on a single proprietary LLM model. Sample images and captions are provided in Table III.
TABLE III: Sample labeled Questions & Answers and image captions from SOTIF-VQA dataset.
Task Image 1 Image 2
Image [Uncaptioned image] [Uncaptioned image]
Close-ended QA Q1: Is the visibility of road signs and pedestrians clear in this image? Q1: How many vehicles can you discern through the windshield?
A1: No A1: Four
Q2: Does the image show a clear view of the traffic lights? Q2: Is visibility reduced in this image due to weather?
A2: Yes A2: Yes
Open-ended QA Q: What specific environmental factor in the image degrades the perception ability of a driver or an autonomous driving system? Q: What are the perception-related SOTIF risks evident in this image due to rain?
A: The low sun angle causing lens flare and reduced visibility is the environmental factor degrading perception. A: The rain creates a wet and reflective road surface, reduced visibility, and potential glare from surrounding vehicle lights, which can degrade the ability of an automated driving system to accurately detect lane markings, other vehicles, and pedestrians.
Q: What technological improvement could help reduce the perception-related SOTIF risk due to direct sunlight? Q: What type of environment subset does this scenario belong to in the context of SOTIF problems?
A: Improving camera dynamic range and employing polarizing filters or advanced image processing algorithms to mitigate glare and enhance visibility could help. A: This scenario belongs to the ’rain’ subset within the environment category, which involves scenarios that degrade the perception ability of automated driving systems due to natural environmental factors.
Image Captioning A driver’s perspective of a busy intersection during sunset, highlighting the challenges of low sun glare in urban driving. Driving in heavy rain, the windshield is covered in water droplets, making visibility of the traffic ahead poor.
Question Generation Agent: The primary objective of this agent is to generate natural, human-like questions based on the input query and captions generated, along with 2D object detection and uncertainty information  [^1] Please refer to supplemental material for object detection and uncertaintyquantification. . Specifically, the question-generating agent is instructed to generate questions with varying difficulties and types. For closed-end questions, the agent randomly selects the question type from Table II. For open-ended questions, the agent focuses on advanced QA like causal inference, including identifying and explaining the SOTIF risk and providing recommended action to prevent the risk, if possible. At the same time, the agent also generates meta data about the question, such as the question type, difficulty level, and expected answer type. Sample questions generated are shown in Table III.
Answer Generation Agent: In the VQA dataset format, each question is answered by ten responses from ten annotators. While this approach ensures accuracy, it is both costly and time-consuming. To streamline this process, the answer generation agent is designed to provide ten answers for each question, given an image. To achieve this, we employ a model aggregation method that utilizes GPT-4o and Claude3-Opus. For each image, the image file and associated questions are fed into the LLMs. Instead of sequentially querying a single model multiple times, the system generates five answers for each question in a single pass through each model. This method is more efficient and cost-effective compared to sending the questions individually. Similarly, the answer generation agent outputs metadata for each answer, including the confidence score and answer type.
Validation and Post-processing Agent: As shown in Figure 2, newly generated answers are then sent to the validation and post-processing agent, which contains three post-processing mechanisms. The first and second mechanisms are to ensure that captions and questions generated are related to the image and perception SOTIF context. The third mechanism is to ensure that answers generated are correct based on the input image and question. If any inconsistencies are found, the agent regenerates the captions, questions or answers as needed. This validation process leverages OpenAI’s GPT4.1 as the LLM backend to ensure high-quality outputs that align with the perception SOTIF context.
III-C Dataset Overview
The multi-agent system employed varying temperature settings during generation to ensure diversity across questions, answers, and captions. Following automated generation, all data entries underwent structural validation and format standardization, with close-ended and open-ended questions automatically categorized using metadata from the generation process. A human review of 595 samples, selected using finite population sampling [^2] More details on the finite population sampling method can be found in S.IV of the supplemental material.., indicated a 3-4% error rate, which is acceptable for automated pipelines. An additional round of manual review addressed edge cases and subtle inconsistencies that automated systems might miss or generate.
The final DriveSOTIF dataset contains 1,114 images, 1,114 captions, and 5,570 question-answer pairs  [^3] Collecting SOTIF-related scenarios is challenging due to their rarity and unpredictability in real-world settings. We are actively collaborating with PeSOTIF authors to expand the dataset with new scenarios and modalities.. The dataset adopts COCO Caption format [27] for image descriptions and VQAv2 format [28] for question-answering tasks. We partition the data using a 90/10 train-test split, as detailed in Table IV.
Table III demonstrates the dataset’s strong alignment with perception-related SOTIF challenges. Quantitatively, caption length averages 18-19 words compared to 10.5 words in COCO Caption, providing substantially richer contextual information essential for understanding complex driving scenarios. Question difficulty maintains excellent balance across easy (39.9%), medium (40.2%), and hard (19.9%) categories. Similarly, question formats split evenly between close-ended (45.7%) and open-ended (54.3%) types, which facilitates assessment of both factual accuracy and reasoning capabilities. Answer types span analysis (36.1%), yes/no and multiple-choice (28.1%), recommendations (17.8%), counting (17.0%), and identification (1.0%) tasks. This diversity enables comprehensive benchmarking of models’ ability to handle both fundamental perception tasks and higher-level causal reasoning and safety recommendations specific to SOTIF.
The generation statistics demonstrate the effectiveness of the multi-agent approach and validate our quality control mechanisms. Caption generation achieved 90.1% first-attempt success, reflecting the straightforward nature of descriptive tasks. Question generation followed with 82.7% success, while answer generation required the most refinement, averaging 1.72 attempts per item due to the complexity of ensuring factual accuracy in analytical responses. Despite this iterative process, 71.0% of answers passed initial validation, with 21.4% requiring single revisions and only 7.6% needing multiple corrections. These results demonstrate that the pipeline consistently produces high-quality, diverse, and contextually appropriate annotations for perception-related SOTIF scenarios.
TABLE IV: Dataset statistics and generation information
Dataset Split Statistics
Train/Valid Test
Number of questions 5,025 555
Number of answers 50,250 5,550
Number of captions 1,005 111
Average question length (words) 11.9 11.8
Average answer length (words) 20.5 21.0
Average caption length (words) 19.22 18.22
Dataset Content Distributions
Difficulty Distribution
 Easy questions 2,224 (39.9%)
 Medium questions 2,243 (40.2%)
 Hard questions 1,113 (19.9%)
Question Type Distribution
 Close-ended questions 2,549 (45.7%)
 Open-ended questions 3,031 (54.3%)
Expected Answer Type Distribution
 Analysis 2,014 (36.1%)
 Yes/No & multiple choice 1,568 (28.1%)
 Recommendation 995 (17.8%)
 Count 947 (17.0%)
 Identification 56 (1.0%)
Generation Information
Average Generation Attempts
 Caption generation 1.16 attempts/item
 Question generation 1.23 attempts/item
 Answer generation 1.72 attempts/item
Success Rates by Attempt Number
 Caption generation 90.1% / 8.3% / 1.6%
 Question generation 82.7% / 13.5% / 3.9%
 Answer generation 71.0% / 21.4% / 7.6%
III-D Evaluation Metrics
To comprehensively evaluate model performance, we focus on two primary aspects: the similarity between model-generated answers and human-annotated ground truth, and the compatibility between generated responses and visual content.
For image captioning, we adopt established metrics including BLEU, METEOR, ROUGE, CIDEr [29], and SPICE [30]. CIDEr computes the average cosine similarity between candidate and reference captions, while SPICE measures precision and recall based on scene graphs. Together, these metrics provide a holistic assessment of caption accuracy, fluency, and contextual relevance.
For close-ended VQA tasks, we use accuracy scores, which measure the percentage of correct answers generated by the model. For open-ended VQA, where no ground truth exists, we follow a rubric-based approach, leveraging an LLM-as-judge [31] to score answers. Specifically, we use GPT-4.1 to evaluate responses based on the question, ground-truth answer, input image, and the model’s output. Each response is scored from 1 to 5 across multiple criteria: relevance, trustworthiness, clarity, and coherence[32]. The evaluation is repeated three times with a temperature of 0.7 to maximize consistency with human judgment.
This evaluation strategy enables nuanced assessment of both factual correctness and the quality of higher-level reasoning in perception-related SOTIF scenarios.
IV Model Benchmarking and Fine-tuning
IV-A Benchmarking
For image captioning benchmarking, we evaluate the Blip [33] and Blip2 models [34] (pre-trained on the COCO Caption dataset [27]) across various model sizes. For VQA tasks, we benchmark a total of Four open-sourced models, including LLaVA 1.5 [35], Qwen2/2.5-VL [36, 37], and InterVL3 [38].
These models are classified by architecture, model size, and quantization method, focusing on two specific variants: 2/3B and 7/8B. The 2/3B variant is particularly suitable for mobile device deployment due to its lightweight design and efficient performance. The 7/8B model is well-suited for on-vehicle computing, where high computational capability is required. By evaluating across this spectrum, we aim to understand trade-offs between accuracy, resource demands, and deployment feasibility for perception-related SOTIF applications in autonomous driving.
IV-B Model Fine-tuning
To enhance the performance of our benchmarked models, we adopt Parameter-Efficient Fine-Tuning (PEFT) techniques such as Low-Rank Adaptation (LoRA) [39] and Quantized LoRA (QLoRA) [40]. LoRA introduces low-rank matrices into each Transformer layer, allowing the model to adapt with far fewer trainable parameters while keeping the original model weights frozen. QLoRA builds on this by incorporating quantization, which further reduces memory requirements and enhances efficiency, all while maintaining performance levels. This makes it easier to adapt LLMs for specialized tasks without needing extensive hardware resources.
For image captioning, we fine-tuned Blip and Blip2 models using the LAVIS framework [41]. For VQA, we utilized LLaVA [35] and LLaMA-Factory [42] frameworks to fine-tune LLaVA 1.5, Qwen2/2.5-VL, and other leading vision-language models. Hyperparameters, random seeds, and training configurations were kept consistent across experiments for fair comparison (details in S.II of the supplemental material).
TABLE V: Performance comparison of baseline models (Base), fine-tuned models (FT), and ground truth (GT) for visual question answering (VQA) and image captioning (IC) tasks.
Task Type Image 1 Image 2
[Uncaptioned image] [Uncaptioned image]
IC Base: a view of a car driving on a wet road a car that is upside down on the side of the road
FT: a driver’s view through a windshield on a rainy day, with red taillights visible ahead a car flips over on a busy road, causing traffic congestion and hindering visibility for drivers on a busy city street
GT: View from a vehicle’s dashboard camera showing a rainy and blurred roadway scene with visible windscreen wiper, a truck ahead, and blurry tail lights An overturned car in the middle of a busy street highlights potential hazards in urban traffic conditions
VQA Q: What weather condition is depicted in the image that could affect perception systems in vehicles? What perception-related SOTIF risk is present due to the vehicle’s unusual position in the scene?
Base: rainy weather, which can affect the perception systems by reducing visibility, making it more difficult for the systems to accurately detect and track objects. camera’s field of view may not capture the full extent of the scene
FT: The image shows heavy rain conditions, which can significantly degrade the performance of perception systems in vehicles by scattering and reflecting light The overturned vehicle poses a risk as it is not commonly expected to be in that position on the road, which can lead to misclassification or delayed recognition by autonomous driving systems.
GT: The image shows heavy rain conditions which can degrade the visibility and accuracy of vehicle perception systems by scattering light and obscuring road markings and other vehicles. The overturned vehicle presents a high SOTIF risk by unexpectedly altering the driving environment, which could lead to inappropriate responses from an autonomous driving system unprepared for such irregular situations.
TABLE VI: Image Captioning Scores
Method Size BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr SPICE
Base FT Base FT Base FT Base FT Base FT Base FT Base FT Base FT
Blip Small 0.348 0.659 0.219 0.464 0.132 0.318 0.082 0.217 0.134 0.208 0.315 0.414 0.138 0.411 0.086 0.143
Large 0.349 0.671 0.227 0.463 0.139 0.312 0.083 0.206 0.137 0.210 0.324 0.399 0.140 0.410 0.091 0.145
[…]
6.7B 0.398 0.723 0.270 0.509 0.169 0.353 0.106 0.248 0.146 0.225 0.338 0.421 0.176 0.536 0.096 0.155
TABLE VII: VQA Scores Comparison: Baseline vs. Fine-tuned Models
Model Size Accuracy (%) Relevance Trustworthiness Clarity Coherence Overall
Base FT Base FT Base FT Base FT Base FT Base FT
LLaVA 1.5 7B 57.51 64.06 3.88 4.01 4.04 4.44 4.69 4.85 4.68 4.89 4.32 4.54
[…]
7B 56.89 62.13 4.24 4.14 4.32 4.51 4.81 4.90 4.79 4.90 4.54 4.61
[…]
7B 54.10 60.12 4.57 4.13 4.66 4.51 4.88 4.90 4.92 4.92 4.76 4.61
[…]
2B 58.01 64.84 4.22 3.92 4.21 4.34 4.74 4.80 4.76 4.82 4.48 4.47
8B 59.12 64.49 4.56 4.12 4.70 4.53 4.94 4.90 4.96 4.92 4.79 4.62
IV-C Results and Analysis
IV-C1 Qualitative Analysis
Table V presents representative examples illustrating how fine-tuning improves model performance on both image captioning and VQA tasks. Baseline captioning models tend to generate short, generic descriptions (e.g., “a car driving on a wet road”) that miss key scene details. After fine-tuning, models produce richer, context-aware captions such as “a driver’s view through a windshield on a rainy day, with red taillights visible ahead,” capturing both environmental context and potential hazards.
For VQA, baseline models typically succeed on simple, factual questions but often fail to provide detailed, situation-specific reasoning. Fine-tuned models demonstrate better understanding of complex queries. For example, when prompted about rain-related perception issues, the fine-tuned model highlights specific challenges like reflections and distortions, offering a deeper analysis of SOTIF risks. In unusual scenarios—such as an overturned vehicle—fine-tuned models identify relevant perception challenges (e.g., blocked sightlines or misclassification risks) that baseline models may overlook.
IV-C2 Quantitative Analysis
Fine-tuned models show consistent and significant improvements across standard metrics. In image captioning, the fine-tuned Blip2 6.7B model achieves ROUGE-L, CIDEr, and SPICE scores of 0.421, 0.536, and 0.155, representing increases of 24.6%, 204.5%, and 61.5% over the baseline, respectively. Similar improvements are observed across BLEU, METEOR, and other benchmarks, indicating enhanced contextual relevance and descriptive accuracy. Notably, the smaller Blip2 2.7B model also performs competitively, highlighting the value of balancing model size and efficiency.
For VQA, larger models generally outperform smaller ones, with the 8B variant of InternVL3 reaching the highest baseline accuracy at 59.1%. Fine-tuning yields substantial accuracy gains for all models: for instance, InternVL3 2B rises from 58.0% to 64.8%. Trustworthiness, clarity, and coherence metrics—scored by LLM-as-judge—also improve, often by more than 0.6 points on a 5-point scale. Of note, fine-tuned Qwen2-VL 2B achieves competitive performance (61.01% accuracy) with only a fraction of the parameters compared to larger models. While accuracy and most quality metrics improve substantially, some metrics such as “relevance” show smaller gains, that even baseline models can identify broad risk factors. However, they lack the detailed, domain-specific knowledge required for precise safety recommendations.
Both qualitative and quantitative results confirm that supervised fine-tuning on the DriveSOTIF dataset significantly enhances model performance for perception-related SOTIF scenarios. The domain-specific dataset enables models to better interpret complex, safety-critical driving situations, supporting more robust and reliable reasoning for autonomous driving applications.
V Experiments
V-A Model Deployment and Real-time Performance Evaluation
MLLMs can be deployed across various environments, including local devices, cloud platforms, and edge computing systems such as multi-access edge computing (MEC). Common deployment frameworks include Huggingface, TensorRT-based solutions [43], and the vLLM framework [44], which was selected for this research due to its efficiency and support for advanced scheduling strategies.
The deployment setup follows a simplified API request-response architecture: input images and text are sent to the deployed model via API calls, and the model returns the corresponding output. For this work, compact, fine-tuned MLLMs were used to ensure fast and resource-efficient inference. These models were merged with their respective adaptive weights and deployed on two platforms: NVIDIA Jetson Orin and a consumer-grade RTX 3090 GPU.
Two inference strategies were considered. Sequential inference processes inputs one at a time as they arrive. This provides deterministic, per-sample latency—an essential requirement for real-time, safety-critical applications like autonomous driving. Continuous batching, by contrast, dynamically aggregates requests and schedules token-level processing across partially completed sequences. Instead of traditional static batching, vLLM uses iteration-level scheduling and PagedAttention to refill GPU slots as soon as they are available. This maximizes GPU utilization and minimizes token-level idle time, resulting in higher throughput without sacrificing latency [44].
V-A1 Sequential Processing
We tested this mode using 111 images from DriveSOTIF, each paired with a prompt for SOTIF risk analysis. Generation time for each image were recorded and aggregated in Table V-A1. GPU VRAM usage was measured via HuggingFace APIs due to vLLM’s memory pre-allocation behavior.
[b]s
TABLE VIII: Real-time performance across hardware platforms and quantization levels.
RTX 3090 Jetson AGX Orin1
Model Size Quant RAM Usage Mean Median Std Mean Median Std
[GB] [s] [s] [s] [s] [s] [s]
Qwen2-VL 2B - 4.74 0.82 0.80 0.11 3.70 3.60 0.27
FP8 2.91 0.74 0.73 0.03 2.93 2.89 0.18
4bit 2.05 0.87 0.86 0.04 3.56 3.49 0.23
[…]
4bit 2.90 1.01 1.00 0.03 5.63 5.61 0.21
[…]
2B - 4.14 1.09 1.07 0.16 4.21 4.14 0.40
  • 1
    We utilized vLLM docker for this testing. Orin is set to MAXN. We encountered system throttling warnings due to over-current.
  • 2
    LMDeploy is used for InternVL3 deployment.
The peak GPU memory usage and processing time per image are recorded. The statistical analysis is performed, and the result is presented in Table V-A1. Based on our results, quantization reduces GPU memory usage while improving real-time performance. Specifically, Qwen2-VL (2B) in 4-bit quantization achieves the lowest memory footprint (2.05 GB) but with a slight latency increase compared to FP8. FP8 quantization yields the best speed on Jetson Orin, reducing inference time by approximately 21% relative to the full-precision model. For larger models like Qwen2.5-VL (3B), 4-bit quantization significantly reduces memory usage from 7.64 GB to 2.90 GB and improves inference speed on both hardware platforms, achieving over 1.3× speedup on Jetson Orin.
Comparing across models, InternVL3-1B consistently exhibits strong performance with the lowest latency and compact memory usage, making it suitable for deployment in resource-constrained environments. However, as the model size increases (e.g., InternVL3-2B), the latency grows accordingly, especially on edge devices.
V-A2 Continuous Processing for high-speed driving
To evaluate continuous inference, we replayed four 30 Hz video streams of high-speed driving scenarios, triggering model inference on every frame. Each frame generates a request to the InternVL3-1B vision-language model, which is deployed on an RTX 3090 with continuous batching enabled  [^4] Please refer to Section V of the supplemental material for results on RTX4090 and duel RTX4090 setup. The system supports multiple requests in parallel, and both queuing and auto-timeout mechanisms are in place: when more requests arrive than the hardware can process immediately, excess requests are placed in a first-in, first-out (FIFO) queue, and any request waiting or processing beyond a set timeout threshold is automatically dropped. We varied the maximum number of in-flight (active) requests—from 1 up to 30 (concurrency)—and, for each setting, recorded the API response times and the number of active requests throughout the experiment (see Figure 3).
At low concurrency levels (1, 3, 5), response times remained stable and low, averaging around 0.5 to 0.7 seconds per frame. As concurrency increased past 10, response times began to rise rapidly, reaching an average of 3.5 seconds per frame at the highest level tested (30). This sharp jump signals that the model and GPU are saturated, leading to longer waits and increased resource contention.
We visualized these dynamics by jointly plotting queue wait times and API response times for each concurrency setting (Figure 4). At low concurrency, the system processes requests quickly once they reach the model, but incoming requests rapidly accumulate in the queue, leading to prolonged wait times and, frequently, discarded requests due to timeouts. As concurrency increases, response times become more variable and overall latency rises, but queue wait times decrease since more requests are processed in parallel. In contrast, with very low concurrency,
This illustrates a fundamental tradeoff between response time and queue waiting time with various concurrency levels. High concurrency leads to longer per-request processing times but keeps queue wait times low. Low concurrency minimizes processing latency for individual requests but causes significant queuing and dropped requests under heavy load. For real-time applications such as autonomous driving, it is essential to maintain concurrency at a moderate level to balance response latency and throughput, ensuring timely and reliable perception without excessive delays or data loss.
Figure 3: Average generation time per image in continuous inference mode.
Figure 4: Tradeoff between API reponse time and queue waiting time in continuous inference mode.
V-B Real World Data Evaluation
For real-world evaluation, as shown in Table V-B, we evaluated the fine-tuned models using data collected in Waterloo, Ontario, Canada, along with proprietary dash-cam footage from China. This evaluation is conducted through a ROS code that queries the MLLM API with random questions from a list, along with image data from collected data in the form of ROSBag or video.
In detail, we used image and video data from the University of Waterloo’s WATonoBus project [45] and the Canadian Adverse Driving Conditions (CADC) dataset [46]. The WATonoBus is an autonomous shuttle operating along a 2.7-kilometer, two-lane University Ring Road that encircles the campus. The CADC dataset was gathered by an autonomous vehicle platform navigating various areas around Waterloo.
The first scenario features a road maintenance crew working on a snowy road inside the university campus with a snowplow in operation. The snowplow, being a rare object, along with the workers and traffic cones, pose challenges for perception systems. The second scenario involves a truck and trailer preparing to merge into the ego vehicle’s lane on a snowy, slippery residential road, where visibility is reduced due to the adverse weather conditions. These two scenarios effectively demonstrate the SOTIF challenges faced by autonomous vehicles in both public roads and university campus environments.
In China, we evaluated the models on two challenging scenarios. The first scenario involved a child hidden in a flattened box lying on the ground, with only a small portion of his foot visible. This scenario is particularly difficult, as the hazard is subtle and easily missed by both human drivers and autonomous systems. In this case, the human driver successfully stopped in time, avoiding an accident. Our aim is to examine whether the fine-tuned MLLM can detect these tiny yet critical visual cues to prevent such accidents. The second scenario took place at night, where the ego vehicle was driving on a wet, reflective road surface due to rain. The glare from the headlights of an oncoming vehicle further complicated the situation, making it difficult to perceive the road and surrounding environment clearly. In this scenario, an electric bicycle crossed the road median unexpectedly and collided with the ego vehicle, resulting in an accident.
TABLE IX: Results of Real-World Evaluation
Waterloo - Snow Plowing, University of Waterloo campus Waterloo - Merging Scenario, city of Waterloo
[Uncaptioned image] [Uncaptioned image]
[…]
A: The risks include potential misidentification of road boundaries due to snow, reduced contrast between the road and its surroundings, and the presence of workers and equipment which could be obscured by snow or poorly illuminated. A: The perception-related SOTIF risk in this image is caused by heavy snowfall, which can reduce visibility and obscure road markings, signs, and other vehicles.
China - Kid hiding in a box China - Strong Vehicle Front Lights, with reflective road
[Uncaptioned image] [Uncaptioned image]
[…]
A: The SOTIF risks include the potential for an autonomous vehicle’s sensors to misinterpret the presence and position of objects on the road, such as the boxes and the bicycle, due to the cluttered and irregular street setup. This could lead to inappropriate vehicle responses or failures to detect critical road users. A: The perception SOTIF risks include failure to detect lane markings, nearby vehicles, or pedestrians due to reduced visibility. The recommended action includes activating hazard lights and slowing down, and possibly engaging in manual driving if the perception system fails.
V-C Results
As shown in Table V-B, in our experiments, fine-tuned MLLMs performed effectively in challenging environments, yet it also highlighted critical areas for enhancement. Notably, in Waterloo’s snowy conditions, the MLLM successfully managed complex scenarios with obscured road markers and temporary obstacles like snowplows and maintenance crews. This performance highlights the model’s robust feature recognition capabilities under adverse weather conditions, proving its potential in dynamic real-world applications.
In the Strong Vehicle Front Lights scenario, despite intense glare from oncoming vehicle headlights, the MLLM successfully identified potential risks and provided reasonable recommended actions. The model recognized that the glare could significantly reduce visibility, increasing the likelihood of missing nearby vehicles or pedestrians. In response, it suggested actions that could directly address the visibility issue and enhance safety, such as slowing down, increasing following distance, or activating hazard lights.
Conversely, the MLLM’s performance in detecting subtle hazards, such as the partially visible child in a box during the China experiments, revealed some limitations. While the model detected unusual patterns associated with the obscured visual cues, it did not accurately capture and interpret the partially or fully obscured child. This highlights a scenario that is challenging even for human drivers. One possible reason is the current capacity of the visual encoders used in the MLLM.
VI Discussion
VI-A Implications of Experimental Results
The experimental findings highlight three key capabilities: accurate risk detection and cognition under adverse condition, causal reasoning linking environmental factors to safety outcomes, and actionable mitigation strategies that address specific SOTIF vulnerabilities. With only 2 GB of GPU memory, the fine-tuned InternVL3-1B model can be deployed to vehicles and replace the API-based vision model for local (near) real-time visual risk assessment. This eliminates privacy concerns, cybersecurity risks, and reliance on internet access.
Pairing this model with a decision-making framework like DriveLLM[45] enables autonomous systems to incorporate SOTIF risks into real-time decision-making. This integration improves safety in complex environments and potentially prevents fatal accidents like the child-hidden-in-the-box scenario, where conventional perception systems may fail to anticipate hidden risks.
Most importantly, deploying a fine-tuned model that supports continuous improvement significantly enhances adaptability. By continuously updating the dataset and applying techniques like SFT and reinforcement learning from human feedback (RLHF) [47], the MLLM can progressively refine its performance and extend to other tasks such as scenario understanding and ODD assessment.
VI-B Domain Transfer
From the labeled response presented in Table III, proprietary LLM models demonstrate the ability to provide accurate and relevant answers to complex, domain-specific questions related to AD perception and SOTIF risks, even without prior specialized training. This highlights their effectiveness in zero-shot and one-shot scenarios, where little or no task-specific data is available.
Moreover, as seen in Tables V and V-B, SFT processes enable baseline models to incorporate domain-specific knowledge. These fine-tuned models perform reasonably well when answering questions about SOTIF challenges for vehicles operating in both Waterloo, Canada, and China. This indicates that the models can effectively transfer learned knowledge across different environmental contexts, adapting to diverse traffic scenarios, weather conditions, and regulatory landscapes.
That said, there are certain limitations to this approach. While domain transfer allows for generalization across different contexts, model performance may be less consistent in environments with extreme conditions or significantly different operational parameters. Although the models adapt well to many contexts, some scenarios may still require additional fine-tuning to maintain accuracy and reliability, especially when facing highly specific regulatory or environmental challenges.
VI-C Hallucination
Hallucination in LLMs refers to the generation of plausible yet factually incorrect or inconsistent responses. It represents a critical challenge for safety-critical applications like autonomous driving. Hallucination can occur at various stages, including pre-training, post-training, SFT, and inference[48]. Within this research, we focus specifically on hallucination suppression during the SFT stage.
During dataset generation, we observed instances of object hallucination, where VLMs would inconsistently report the presence of objects across multiple queries. For example, when asked whether a specific object was present in an image, one out of three attempts indicated ”yes,” while the remaining attempts responded ”no.” This inconsistency exemplifies fact-conflicting hallucination, a critical reliability issue in VLMs where models generate contradictory responses to identical queries [48].
To address this challenge, we implemented a multi-stage hallucination suppression strategy during the SFT process. Our approach includes: (1) a post-processing validation step using an independent LLM to cross-verify generated responses against input images, (2) consistency filtering that removes outputs failing validation checks, and (3) iterative correction of inconsistent annotations. This targeted intervention significantly reduces hallucinated content entering the fine-tuning pipeline, ensuring our dataset maintains grounding in factual, image-based annotations.
However, inference-time hallucination remains a persistent challenge, particularly due to inherent knowledge gaps and biases in pre-trained model backbones. Existing mitigation strategies present distinct trade-offs for real-time applications. Retrieval-augmented generation [49], while effective at grounding responses in factual content, introduces additional inference latency that may be incompatible with the sub-second response requirements of autonomous driving systems. Human feedback mechanisms [50], though widely adopted in industry and demonstrably effective at improving model reliability, face scalability constraints when applied to high-frequency inference scenarios typical in autonomous vehicles. At the moment, one feasible direction for addressing inference-time hallucination is grounding model response through a human-in-loop semi-auto feedback mechanism. Other methods, like latent space steering [51], can also be explored to reduce hallucination while maintaining real-time performance.
VI-D Future of LLM in SOTIF
As current research on LLMs in SOTIF applications remains limited, this section outlines potential future uses of both text-based and multimodal LLMs. First, a universal description system for SOTIF scenarios can be developed to provide standardized descriptions. Building on this foundation, custom-trained LLMs can generate JSON-based representations for scenario detection, communication, and risk identification. This standardization would make it easier to share and compare SOTIF scenarios across different research groups and industry applications.
These standardized descriptions would then support better testing approaches. LLMs can efficiently generate SOTIF test cases following systematic testing plans. By learning from test failures, these systems can improve subsequent tests, reducing the total number of tests needed. LLMs can also create highly targeted test cases for specific SOTIF risks, focusing on edge cases and rare conditions that are most challenging and potentially dangerous.
The testing capabilities naturally extend to evaluation tasks. Multimodal LLMs can assist in evaluating existing rule-based or learning-based SOTIF methods. By providing advanced scene understanding, MLLMs can refine manually defined evaluation algorithms and cross-check results from conventional systems. This ensures SOTIF risks are assessed from multiple angles. MLLMs can also select specific scenarios of interest based on input prompts.
Finally, multimodal LLMs can integrate data from multiple sensor types, such as LiDAR, radar, and bird’s-eye-view images. By encoding these inputs into feature embeddings alongside textual prompts, LLMs could directly assess SOTIF risks at the feature level. This approach would provide more detailed risk understanding, reduce reliance on processed detection results, and potentially identify hazards earlier in the perception pipeline.
VII Conclusion and Future Work
Our research demonstrates that SFT techniques applied to MLLMs using the DriveSOTIF dataset can enhance perception-related SOTIF risk detection in autonomous driving systems. By introducing a small domain-specific dataset intended for model fine-tuning, we were able to evaluate strengths and limitations of MLLMs in perception-related SOTIF problems. Benchmark results indicate that our fine-tuned MLLMs outperform baseline models both qualitatively and quantitatively.
These improvements enhance risk identification and situational awareness, even under challenging conditions such as adverse weather and occlusion-heavy scenarios. This is confirmed by experiments on data collected from both China and Canada. Real-world performance testing further confirms that the deployed models achieve an average inference time of 0.59 second per image, making them suitable for low- to moderate-speed applications like urban navigation, automated parking, and traffic jam assistance.
More importantly, by integrating fine-tuned MLLMs with LLM-based decision-making frameworks, autonomous vehicles can incorporate perception risks into real-time operations, ultimately improving road safety. Moreover, the ability of MLLMs to interpret complex scenarios through natural language reasoning establishes concrete foundations for transparent safety assessment through natural language explanations of risk factors and mitigation strategies. Lastly, the on-device and MEC-based deployment of these lightweight models can lead to a privacy-preserving, onboard-cloud hybrid LLM deployment for next-generation autonomous vehicles.
For future work, we plan to improve real-time performance for high-speed safety-critical applications, and minimize hallucinations during inference stage. This will require interdisciplinary collaboration spanning natural language processing and autonomous driving to realize the full potential of MLLMs in safer autonomous driving systems.
Acknowledgment
The authors would like to acknowledge the financial support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Vector Institute. We also acknowledge the research grant provided by OpenAI.